OpenStudy (wesley):

Factor each polynominal 44a^2+11a Can anyone show me how to work this?

OpenStudy (shadowfiend):

Well, the first thing in these cases is always to see that a * a = a^2, so you have: \[\begin{align} 44a^2 &+ 11a\\ 44a\cdot a &+ 11a \end{align}\] In that case, you can factor an \(a\) out: \[a(44a + 11)\] Does that make sense?

OpenStudy (wesley):

Yes

OpenStudy (anonymous):

You can also factor out an 11: \[11a(4a+1)\]